OpenStudy (anonymous):

Which equation can be used to solve the problem? Juanita hiked 5 1/4 miles on Saturday. That was 3/4 as far as Laura hiked. How far did Laura hike? A. m = 3/4 x 5 1/4 B. 3/4m = 5 1/4 C. 3/4 / m = 5 1/4 D. m - 3/4 = 5 1/4

OpenStudy (anonymous):

I'm thinking B or D

Miracrown (miracrown):

What makes you think that its either B or D?

Miracrown (miracrown):

Divide them?

Miracrown (miracrown):

\[\frac{ 21 }{ 4 } \space \div \frac{ 3 }{ 4} \space = \]

OpenStudy (anonymous):

7/1

Miracrown (miracrown):

\[5 \frac{ 1 }{ 4 } \space \div \frac{ 3 }{ 4 } \space = 5\frac{ 1 }{4 } \space \times \frac{ 4 }{ 3 }\]

Miracrown (miracrown):

\[\frac{ 3 }{ 4 } m \space = 5 \frac{ 1 }{ 4 }\]

Miracrown (miracrown):

m represents the miles that Laura hiked. so 3/4*m=5 1/4 Make sense? :)

OpenStudy (anonymous):

yes, I knew it was B or D because when I did the math on all they were the only two that were more than 5 1/4.

Miracrown (miracrown):

Yeah, that's it! Good job :)

OpenStudy (anonymous):

Thanks for you help :)